OpenStudy (anonymous):

I need some help graphing this. (2x^2+2x-4)/(x^2+x). I have to show the slant asymptote and vertical asymptotes. I found that the zeros are 1 and -2, the horizontal asymptote is y=2. What I don't know is the vertical asymptotes. I think the are at x=-1 and 0 because that is what makes the denominator=0. And I think I know that as the limit as x approaches -1 from both sides, it appoaches negative infinity, so the graph goes down. I am getting confused on whether the x=0 is also a VA and if so how to deal with it.

OpenStudy (ranga):

\(\Large \frac{2x^2+2x-4}{x^2+x} = \frac{2(x^2+x-2)}{x(x+1)} = \frac{2((x+2)(x-1)}{x(x+1)}\) Zeros at x = -2 and x = 1. Vertical asymptotes at x = 0 and x = -1.

OpenStudy (anonymous):

Thanks! I think I had it... just got a little mixed up when I did one of the limits. Makes sense now though...

OpenStudy (ranga):

No slant asymptotes here because the degree of the numerator is NOT one more than the denominator.